Deep Thought Granola is 25% nuts and dried fruit. Oat Dream Granola is 10% nuts and dried fruit. How much of Deep Thought and how much of Oat Dream should be mixed to form a 20-lb batch of granola that is 19% nuts and dried fruit?
*I would like to know how to solve the problem step-by-step please. Thank you in advance.

Answer by lwsshak3(11628)

Deep Thought Granola is 25% nuts and dried fruit. Oat Dream Granola is 10% nuts and dried fruit. How much of Deep Thought and how much of Oat Dream should be mixed to form a 20-lb batch of granola that is 19% nuts and dried fruit?
**
let x=lbs of Deep Thought Granola to be mixed
20-x=lbs of Oat Dream Granola to be mixed
..
25%x+10%(20-x)=19%*20
.25x+2-.1x=3.8
.15x=1.8
x=1.8/.15=12
20-x=8
ans:
lbs of Deep Thought Granola to be mixed=12
lbs of Oat Dream Granola to be mixed=8